# Characterizing Spin-Orbit Torques by Tensorial Spin Hall Magnetoresistance

### v1

@@ Acknowledgements.
The authors are grateful to Luqiao Liu, Simranjeet Singh, Andrew Kent, and Haoyu Liu for helpful discussions. This work was initially supported by the Air Force Office of Scientific Research under Grant No. FA9550-19-1-0307, and completed with the support of the National Science Foundation under Award No. DMR-2339315.
Appendix A End Matter
@@
×$\qquad\times\left[\tilde{g}_{R}\epsilon_{jkl}\epsilon_{lab}\hat{m}_{k}\hat{m}_{a}\mu_{b}^{0}+\tilde{g}_{I}\epsilon_{jkl}\hat{m}_{k}\mu_{l}^{0})\right],$ (3)
@@
×$\times\frac{\sinh[(z-t)/\lambda]}{\sinh(t/\lambda)}\left[\tilde{g}_{R}(\chi_{zjx}\chi^{\prime}_{xzk}\hat{m}_{j}\hat{m}_{k}-\chi_{zkx}\chi^{\prime}_{xzk})\right.$
@@
×$\times\frac{\sinh[(z-t)/\lambda]}{\sinh(t/\lambda)}\left[\tilde{g}_{R}(\chi_{zjx}\chi^{\prime}_{yzk}\hat{m}_{j}\hat{m}_{k}-\chi_{zkx}\chi^{\prime}_{yzk})\right.$
@@
According to Eq. (3a), the longitudinal resistivity can be expressed as $\rho_{xx}(\tilde{\bm{m}}_{\perp})=\rho_{0}(\hat{\bm{h}}\cdot\hat{\bm{s}})\rm{Re}(\tilde{\bm{m}}_{\perp}\cdot\hat{\bm{s}})$, where $\rho_{0}$ is an overall factor absorbing $\sigma$, $\tilde{g}_{R}$, $t$, $\eta$ and $|\bm{s}|$. Substituting Eq. (6) into $\rho_{xx}(\tilde{\bm{m}}_{\perp})$ and using the near-resonance approximation $(H^{2}-\omega_{0}^{2})^{2}\approx 4\omega_{0}^{2}(H-\omega_{0})^{2}$, we arrive at Eqs. (6) and (7).
Derivation of second harmonic Hall voltage—In the low-frequency regime, the magnetizaion vector $\hat{\bm{m}}$ remains in quasi-equilibrium, so all time derivatives in Eq. (5) can be set to zero, yielding
@@
$\frac{d\bm{\theta}}{d\bm{H}_{\rm{F}}}\!=\!\left[\left(\mathcal{M}\frac{d\hat{\bm{m}}}{d\bm{\theta}}\right)^{T}\left(\mathcal{M}\frac{d\hat{\bm{m}}}{d\bm{\theta}}\right)\right]^{-1}\!\!\left(\mathcal{M}\frac{d\hat{\bm{m}}}{d\bm{\theta}}\right)^{T}\!(\hat{\bm{m}}\times\mathcal{I}),$ (9)

### v3

@@ Acknowledgements.
The authors are grateful to Luqiao Liu, Simranjeet Singh, Andrew Kent, and Haoyu Liu for helpful discussions. This work was initially supported by the Air Force Office of Scientific Research under Grant No. FA9550-19-1-0307, and completed with the support of the National Science Foundation under Award No. DMR-2339315. All raw data corresponding to the findings in this paper is available upon reasonable requests.
Appendix A Supporting Information
@@
$\qquad\times\left[\tilde{g}_{R}\epsilon_{jkl}\epsilon_{lab}\hat{m}_{k}\hat{m}_{a}\mu_{b}^{0}+\tilde{g}_{I}\epsilon_{jkl}\hat{m}_{k}\mu_{l}^{0})\right],$ (3)
@@
$\times\frac{\sinh[(z-t)/\lambda]}{\sinh(t/\lambda)}\left[\tilde{g}_{R}(\chi_{zjx}\chi^{\prime}_{xzk}\hat{m}_{j}\hat{m}_{k}-\chi_{zkx}\chi^{\prime}_{xzk})\right$.
@@
$\times\frac{\sinh[(z-t)/\lambda]}{\sinh(t/\lambda)}\left[\tilde{g}_{R}(\chi_{zjx}\chi^{\prime}_{yzk}\hat{m}_{j}\hat{m}_{k}-\chi_{zkx}\chi^{\prime}_{yzk})\right$.
@@
[…]
Derivation of 2nd harmonic Hall voltage—In the low-frequency regime, the magnetizaion vector $\hat{\bm{m}}$ remains in quasi-equilibrium, so all time derivatives in Eq. (5) can be set to zero, yielding
@@
[…]